OpenStudy (coolgirljenn):

Briana guesses at all 10 true/false questions on her history test. Find each probability. P(at most half correct) a.) 105/512 b.)319/512

OpenStudy (mww):

Use niomial probability formula. If there are only two possible outcomes, P and Q then there probabilities p and q add to one. And (p+q)^n = 1 The probability of an event such that outcome P occurring r times in n trials is then \[\left(\begin{matrix}n \\ r\end{matrix}\right) p^r q^{n-r}\] P(at most half correct) is the same as saying you can have up to 5 correct, so you include the following: P(none correct) + P(1 correct) + P(2 correct)+...+P(5 correct) The sum of these probabilities gives you your answer.